# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 7.2. Coordinate changes
$E_{L_{\varepsilon}}(\phi)=\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos\phi\,dx$
and $\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})=-\partial_{x}^{2}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sin m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}$. On the right-hand side of (7.10), the first two terms and the last term are error terms, whereas the term $\langle v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})v)\rangle$ defines the covariance operator of the new Gaussian measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}$. Combining (7.8) and (7.10) yields
$Z[F]$ $=\int\cdots\int_{U_{\varepsilon}}\exp\Big\{-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle\Big\}$
$\hphantom{XXXXX}\cdot\exp\Big\{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\Big\}\exp\Big\{-\frac{1}{2}\langle v,\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})v\rangle\Big\}$
$\hphantom{XXXXX}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,\mu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{Q}).$ (7.11)
We first study the first term $\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})$ in (7.11). From Lemma 3.4 and (4.6), we have $E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})=|Q|E_{\textup{kink}}+O(e^{-cd_{\varepsilon}})$, uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$. This implies
$\exp\Big\{-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\Big\}$ $=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}-\frac{1}{\varepsilon}O(e^{-cd_{\varepsilon}})\Big\}$
$=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}(1+O(\varepsilon^{0+}))\Big\},$ (7.12)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}=(1+\eta)\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
We now study the second term $\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle$ in (7.11). By Lemma 3.6 and the condition $\|\sqrt{\varepsilon}v\|_{L^{2}}\leq\delta_{\varepsilon}=\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}$, we have
$\bigg|\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}vdx\bigg|\leq\frac{1}{\varepsilon}\|\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\|_{L^{2}}\|\sqrt{\varepsilon}v\|_{L^{2}}\lesssim\frac{1}{\varepsilon}e^{-(1-\frac{\eta}{2})d_{\varepsilon}}\delta_{\varepsilon}=O(\varepsilon^{0+}),$ (7.13)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}=(1+\eta)\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$. This implies that
$\exp\Big\{-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle\Big\}=1+O(\varepsilon^{0+})$ (7.14)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$.
We now study the second-order term $\frac{1}{2}\langle v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})v\rangle$ in (7.10), which defines a new base Gaussian measure on the normal space. Note that
$-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{Q}\text{sech}^{2}(\cdot-\xi_{j})+O(e^{-cd_{\varepsilon}})=\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}+O(e^{-cd_{\varepsilon}}),$
where the error is independent of $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$. Since $v\in V_{\xi_{1},\dots,\xi_{Q}}$ satisfies the orthogonality conditions, from Lemma 5.4, we can define the Gaussian measure
$d\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}=Z_{\xi_{1},\dots,\xi_{Q}}^{-1}\exp\Big\{-\frac{1}{2}\langle v,\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}v\rangle\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}dv(x),$ (7.15)
Finally, we study the surface measure $d\sigma(\xi_{1},\dots,\xi_{Q})$ on the multi-soliton manifold. Indeed, since the tangent vectors $\{\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}\}_{j=1}^{Q}$ are almost orthogonal, with an error of order $e^{-cd_{\varepsilon}}$, the Jacobian matrix
$M(\xi_{1},\dots,\xi_{Q})=(\langle\partial_{\xi_{i}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}},\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}\rangle)_{1\leq i,j\leq Q}$
is nearly diagonal, with diagonal entries given by $\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}$, which are independent of $\xi_{1},\dots,\xi_{Q}$. Therefore, we have
$\textup{det}(M(\xi_{1},\dots,\xi_{Q}))=Q\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd_{\varepsilon}})),$ (7.16)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$.
By using (7.8), (7.11), (7.12), (7.14), (7.15), and (7.16), we can take out terms independent of tangential modes $\xi_{1},\dots,\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$ as follows:
$Z_{\varepsilon}[F]$ $=J_{\varepsilon}\int\cdots\int_{U_{\varepsilon}}F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,d\nu^{\perp}_{m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}}(v)\,d\xi_{Q}\dots d\xi_{1},$
where
$J_{\varepsilon}=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}\Big\}(Z_{0})^{Q}\cdot Q\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd_{\varepsilon}}))(1+O(\varepsilon^{0+})).$
@@ 8. Analysis of the Gaussian measure with Schrödinger operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$ (8.1)
@@ 8.1. Correlation decay
In this subsection, we study the decay of correlations for the Gaussian measure with covariance operator $C_{\xi_{1},\dots,\xi_{k}}$ in (8.1). Notice that the operator
$-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
is invertible on the normal space $V_{\xi_{1},\dots,\xi_{k}}$. Therefore, we can take its inverse to study the Green function
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y):=\langle\delta_{x},C_{\xi_{1},\dots,\xi_{k}}\delta_{y}\rangle=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big]$ (8.2)
@@
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)=\big(\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big)(x,y)$
with Dirichlet boundary conditions on $L_{\varepsilon},L_{\varepsilon}]$. Note that Dirichlet boundary conditions suppress the variance $(-\partial_{x}^{2}+1)^{-1}(x,x)\approx 0$ near the edges, eliminating randomness there. In particular, the influence of the boundary diminishes exponentially fast as one moves into the interior, and the field in the bulk behaves almost like the infinite-volume Ornstein–Uhlenbeck field. The following proposition shows that the Green function $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ exhibits correlation decay away from the boundary points $L_{\varepsilon}$, $L_{\varepsilon}$.
@@
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
@@
We expand the covariance operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$. Recall that the projection operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j}$ (5.12) onto the normal space $V_{\xi_{1},\dots,\xi_{k}}$ is
$\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ $=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j},$
where each $\mathbf{P}_{j}$ denotes the projection onto the corresponding tangential direction
$\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}.$ (8.3)
Here $t_{1},\dots,t_{k}$ are obtained by applying the Gram–Schmidt orthonormalization to the tangent vectors $\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$, $j=1,\dots,k$. By expanding the projection operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$,
@@
$\hphantom{X}-\sum_{i=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}$ (8.4)
@@
$G_{\text{OU}}^{D}(x,y)=\frac{\text{sinh}(L_{\varepsilon}-\max\{x,y\})\cdot\text{sinh}(L_{\varepsilon}+\min\{x,y\})}{\text{sinh}(2L_{\varepsilon})}$ (8.5)
@@
$\big|G_{\text{OU}}^{D}(x,y)-\frac{1}{2}e^{-|x-y|}\big|\lesssim e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})}\lesssim e^{-2|\log\varepsilon|}.$ (8.6)
@@
$\big\langle\delta_{x},\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}\delta_{y}\big\rangle=t_{i}(x)t_{j}(y)\big\langle t_{i},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle.$ (8.7)
@@
$\big\langle\delta_{x},(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}\delta_{y}\big\rangle$ $=t_{j}(y)\big\langle\delta_{x},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle=t_{j}(y)\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz,$ (8.9)
@@
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}.$ (8.10)
@@
$|\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}|\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-|x-z|}e^{-|z-\xi_{j}|}dz\lesssim e^{-c|x-\xi_{j}|}.$ (8.11)
@@
$\lesssim e^{|\log\varepsilon|}e^{-2(L_{\varepsilon}-\max\{|x|,|\xi_{j}|\})}+e^{-c|\log\varepsilon|}$
$\lesssim e^{-|\log\varepsilon|}+e^{-c|\log\varepsilon|},$ (8.12)
@@
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz\bigg|\lesssim e^{-c|x-\xi_{j}|}$ (8.13)
@@
$\big|\big\langle\delta_{x},\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\delta_{y}\big\rangle\big|\lesssim e^{-c|x-y|}.$ (8.15)
@@
We now prove part (ii) of Lemma 8.2 (diagonal regime). Using the closed form (8.5), we have
$0\leq G^{D}_{\text{OU}}(x,x)\leq\frac{1}{2}$
for all $x\in[-L_{\varepsilon},L_{\varepsilon}]$. Since $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ is an orthogonal projection, so $0\leq\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\leq\operatorname{Id}$. Hence, adding the projection cannot increase the diagonal
$0\leq G_{\textup{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,x)\leq G^{D}_{\text{OU}}(x,x)\leq\frac{1}{2}$
for all $x\in[-L_{\varepsilon},L_{\varepsilon}]$, uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$. This completes the proof of Lemma 8.2.
We are now ready to prove the main part of this subsection. Recall the definition of the Green function corresponding to the covariance operator $\mathcal{C}_{\xi_{1},\dots,\xi_{k}}$ in (8.1)
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y):=\langle\delta_{x},C_{\xi_{1},\dots,\xi_{k}}\delta_{y}\rangle=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big].$
The following proposition shows that the Green function $G_{\xi_{1},\dots,\xi_{k}}$ exhibits correlation decay away from the boundary points $L_{\varepsilon}$, $L_{\varepsilon}$.
@@
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
@@
Using the resolvent identity, we view the Schrödinger operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$
as a perturbation of the Ornstein–Uhlenbeck operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ as follows
@@
$=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)+(G^{D}_{\xi_{1},\dots,\xi_{k}}WG^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}})(x,y),$ (8.17)
@@
$\sum_{j=1}^{k}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
$=\sum_{j=1}^{k}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz,$ (8.18)
where $W_{j}(z)=\text{sech}^{2}(z-\xi_{j})$. Note that the equality (8.18) holds after summing over $j$, although the individual components do not commute. Indeed, by Lemma 8.4, we have
$\big\|\big(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}\big)-\big(G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\big)\big\|_{L^{2}\to L^{2}}\lesssim e^{-c\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|},$ (8.19)
where $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
Using Lemma 8.2, under the bulk condition $|x|,|y|\leq L_{\varepsilon}-|\log\varepsilon|$ we write
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)$ $=e^{-|x-z|}+O(e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})})$
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)$ $=e^{-|y-z|}+O(e^{-2(L_{\varepsilon}-\max\{|y|,|z|\})}).$ (8.20)
@@
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\lambda|x-z|}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz+e^{-c|\log\varepsilon|}$ (8.21)
@@
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\lambda|y-z|}dz+e^{-c|\log\varepsilon|}.$ (8.22)

### v2

@@ 7. Change of coordinates
$J_{\varepsilon}=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}\Big\}(Z_{0})^{Q}\cdot Q\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd_{\varepsilon}}))(1+O(\varepsilon^{\frac{1}{2}-})).$
@@ 8. Analysis of the Gaussian measure with Schrödinger operator
(8.1) $C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$
@@ 8.1. Correlation decay
In this subsection, we study the decay of correlations for the Gaussian measure with covariance operator $C_{\xi_{1},\dots,\xi_{k}}$ in (8.1). Notice that the operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$, defined in (5.5), is invertible on the normal space $V_{\xi_{1},\dots,\xi_{k}}$. Therefore, we can take its inverse to study the Green function
(8.2) $G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y):=\langle\delta_{x},C_{\xi_{1},\dots,\xi_{k}}\delta_{y}\rangle=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big]$
@@
(8.3) $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)=\big(\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big)(x,y)$
with Dirichlet boundary conditions on $L_{\varepsilon},L_{\varepsilon}]$. Note that Dirichlet boundary conditions suppress the variance $(-\partial_{x}^{2}+1)^{-1}(x,x)\approx 0$ near the edges, eliminating randomness there. In particular, the influence of the boundary diminishes exponentially fast as one moves into the interior, and the field in the bulk behaves almost like the infinite-volume Ornstein–Uhlenbeck field. The following result shows that the Green function $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ exhibits correlation decay away from the boundary points $L_{\varepsilon}$, $L_{\varepsilon}$.
@@
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$.
@@
We expand the covariance operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$. Recall the projection operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j}$ (5.13) onto the normal space $V_{\xi_{1},\dots,\xi_{k}}$. By expanding
@@
(8.4) $\hphantom{X}-\sum_{i=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1},$
@@
(8.5) $G_{\text{OU}}^{D}(x,y)=\frac{\text{sinh}(L_{\varepsilon}-\max\{x,y\})\cdot\text{sinh}(L_{\varepsilon}+\min\{x,y\})}{\text{sinh}(2L_{\varepsilon})}$
@@
(8.6) $\big|G_{\text{OU}}^{D}(x,y)-\frac{1}{2}e^{-|x-y|}\big|\lesssim e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})}\lesssim e^{-2|\log\varepsilon|}.$
@@
(8.7) $\big\langle\delta_{x},\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}\delta_{y}\big\rangle=t_{i}(x)t_{j}(y)\big\langle t_{i},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle.$
@@
(8.9) $\big\langle\delta_{x},(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}\delta_{y}\big\rangle$ $=t_{j}(y)\big\langle\delta_{x},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle=t_{j}(y)\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz,$
@@
(8.10) $=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}.$
@@
(8.11) $|\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}|\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-|x-z|}e^{-|z-\xi_{j}|}dz\lesssim e^{-c|x-\xi_{j}|}.$
@@
(8.12) $\lesssim e^{|\log\varepsilon|}e^{-2(L_{\varepsilon}-\max\{|x|,|\xi_{j}|\})}+e^{-c|\log\varepsilon|}\lesssim e^{-|\log\varepsilon|},$
@@
(8.13) $\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz\bigg|\lesssim e^{-c|x-\xi_{j}|}$
@@
(8.15) $\big|\big\langle\delta_{x},\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\delta_{y}\big\rangle\big|\lesssim e^{-c|x-y|}.$
@@
We now prove part (ii) of Lemma 8.2 (diagonal regime). Using the closed form (8.5), we have $0\leq G^{D}_{\text{OU}}(x,x)\leq\frac{1}{2}$ for all $x\in[-L_{\varepsilon},L_{\varepsilon}]$. Since $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ is an orthogonal projection, so $0\leq\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\leq\operatorname{Id}$. Hence, adding the projection cannot increase the diagonal $0\leq G_{\textup{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,x)\leq G^{D}_{\text{OU}}(x,x)\leq\frac{1}{2}$ for all $x\in[-L_{\varepsilon},L_{\varepsilon}]$, uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$. This completes the proof.
We are now ready to prove the main part of this subsection. Recall the Green function $G^{D}_{\xi_{1},\dots,\xi_{k}}$ (8.2) for the covariance operator $\mathcal{C}_{\xi_{1},\dots,\xi_{k}}$ in (8.1). The following proposition shows that the Green function exhibits correlation decay away from the boundary points $L_{\varepsilon}$, $L_{\varepsilon}$.
@@
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$.
@@
Using the resolvent identity, we view the Schrödinger operator $C_{\xi_{1},\dots,\xi_{k}}$, defined in (8.1), as a perturbation of the Ornstein–Uhlenbeck operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ as follows
@@
(8.17) $=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)+(G^{D}_{\xi_{1},\dots,\xi_{k}}WG^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}})(x,y),$
@@
(8.18) $\sum_{j=1}^{k}(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}})$ $=\sum_{j=1}^{k}(G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}),$
where $W_{j}(z)=-2\text{sech}^{2}(z-\xi_{j})$. Note that the equality (8.18) holds after summing over $j$, although the individual components do not commute. Indeed, by Lemma 8.4 below, we have
(8.19) $\big\|\big(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}\big)-\big(G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\big)\big\|_{L^{2}\to L^{2}}\lesssim e^{-c\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|},$
which will be used below. Using Lemma 8.2, we write
(8.20) $G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x_{1},x_{2})$ $=e^{-|x_{1}-x_{2}|}+O(e^{-2(L_{\varepsilon}-\max\{|x_{1}|,|x_{2}|\})}).$
@@
(8.21) $\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$ $\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\lambda|x-z|}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz+e^{-c|\log\varepsilon|}$
@@
(8.22) $\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\lambda|y-z|}dz+e^{-c|\log\varepsilon|}.$